OpenStudy (anonymous):

Solve this quadratic equation using the quadratic formula. x2 + 8x - 5 = 0

pooja195 (pooja195):

List your A,B,C values keeping this in mind \[\huge~\rm~ax^2+bx+c=0\]

OpenStudy (anonymous):

\[\frac{ -8\pm \sqrt{(8)^{2}-4(1)(5)} }{ 2(1) }\]

jigglypuff314 (jigglypuff314):

don't forget about the minus sign (negative) in front of that 5 :)

OpenStudy (anonymous):

thanks

OpenStudy (greencat):

Perfect @gizmo25 Now simplify. \[\frac{ -8 \pm \sqrt{64-20} }{ 2 }\]\[\frac{ -8 \pm \sqrt{44} }{ 2 }\]Oops; I was about to solve it all for you. Can you work it out from here? Hint: Pull out what you can from the square root.

jigglypuff314 (jigglypuff314):

@GreenCat note that I just mentioned the negative so it should be a 64 + 20 instead :)

OpenStudy (greencat):

oh yes; it is (-5). Sorry.

OpenStudy (greencat):

I was working from Gizmo's equation. @gizmo25 Remember to include the negative in "-5" there.

OpenStudy (greencat):

\[\frac{ -8 \pm \sqrt{64+20} }{ 2 }\]\[\frac{ -8 \pm \sqrt{84} }{ 2 }\]Here You go.

OpenStudy (anonymous):

thanks :)